OpenStudy (anonymous):

Solve for x and check. Enter only a number. 3^(2x) = 27

OpenStudy (anonymous):

3^2x = 3^3 can u finish it now?

OpenStudy (anonymous):

what do i do with the threes

OpenStudy (anonymous):

3^3 is only equal to 3^3. therefore for 3^2x = 3^3 then 2x must be equal to 3. thus 2x = 3 x = 3/2

OpenStudy (anonymous):

are you positive

OpenStudy (anonymous):

I'll check @tomtom777

OpenStudy (anonymous):

ok @hedyeh99

OpenStudy (anonymous):

Yes he is correct

OpenStudy (anonymous):

:P dont dought me

OpenStudy (anonymous):

doubt*

OpenStudy (anonymous):

lol sorry